One number is four times another number. The largest number is also 87 more than the smaller numer. Find the numbers.
Choices for the smaller number: 29, 30, 33

Answer by stanbon(75887)

One number is four times another number. The largest number is also 87 more than the smaller numer. Find the numbers.
Choices for the smaller number: 29, 30, 33
----------------------
EQUATIONS:
x = 4y
x-y =87
-----------
Solve by substitution:
4y-y=87
3y = 87
y = 87/3
y = 29 (smaller number)
-----------
x = 4y=4*29 = 116 (larger number)
========================
Cheers,
Stan H.